Solve the following equations
(i) 3- 4x2 = -85
(ii) -5x2 = -500

Answer by algebrahouse.com(1659)

"Solve the following equations
(i) 3- 4x2 = -85
(ii) -5x2 = -500"

3 - 4x² = -85
-4x² = -88 {subtracted 3 from both sides}
x² = 22 {divided both sides by -4}
x = ±√22 {took square root of both sides}

-5x² = -500
x² = 100 {divided both sides by -5}
x = ±10 {took square root of both sides}
www.algebrahouse.com